On graph paper, graph the two lines associated with the equations in this system of equations.
3x + 2y = 13
2x + 3y = 12
Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.

Answer by TimothyLamb(4379)

3x + 2y = 13
2x + 3y = 12
---
put the system of linear equations into standard form
---
3x + 2y = 13
2x + 3y = 12
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3
y = 2
(3,2)
---
to graph them use the slope and y-intercept of each line:
---
3x + 2y = 13
2y = -3x + 13
y = (-3/2)x + 13/2
slope = -3/2
y-intercept = (0,13/2)
---
2x + 3y = 12
3y = -2x + 12
y = (-2/3)x + 12/3
y = (-2/3)x + 4
slope = -2/3
y-intercept = (0,4)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php